You are here: Foswiki>PartnerObserving Web>ItalyQueue2013>ObsInaf20130219 (31 Mar 2013, EleanoraSani)Edit Attach

INAF queue observing, 2013 February 18-19

Observer: E. Sani, M. Fumana
Telescope Operator: S. Allanson
Support Astronomer: D. Thompson
Instrument: LUCI - LBC blue only

Summary

During the late evening the sky is covered by very thick clouds. We execute LUCI calibrations
Problem with the red side door of the dome. Mike tries to open it manually from level 10
It is not possible to fix the problem, we observe with the blue-side alone

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci_DDT HH34 FeII ~24min Interrupted: target close to wind dir
Zamorani_A1 Hip33836 HKspec    
Zamorani_A1 zCgr32 HKspec 2.5h Interrupted: guide star lost
LaFranca_A9 Hip061493 200HK/zJspec    
LaFranca_A9 NGC4338 200HK/zJspec 45min COMPLITED
Grazian_30A GOODS-N U-special 1h12min Only blue channel
Standard SA109 U-special    

Details

00:03 We start LUCI calibrations
  • ./DoDarks.txt
    • 60s 0005-0014
    • 120s 0015-0019
    • 24s 0020-0024
    • 600s 0025-0029
    • 350s 0030-0034
    • 15s 0035-0039
    • 10s 0040-0049
    • 12s 0050-0059
    • 5s 0060-0069
    • 4s 0070-0079

* Since the guiding star of the scientific filed Antoniucci_DDT is close to the recommended limit (R=15.9) and we have heavy thick clouds, we decide to change it with a brighter one (R=14.3). To do this, we had to change the PA from 0.0 to 90.0 and the order of the absolute offsets to prevent the guider from loosing the star.

02:08 Clouds are less thick now. We start observe

02:09 DOME OPENING
  • Problem with the red side door of the dome.

03:22 Since the problem is still not fixed and the day crew is required to solve it. We decide to stay all night long with LUCI and use only the blue side. The TO starts to collimate.

03:31 executeLUCIscript HH34_ima_feII.img 0080
  • since the new guiding star is a double star, there are jumps for guiding. The sky conditions are now better and we switch back to the original guiding settings.

03:37 executeLUCIscript HH34_ima_feII.img 0081-0092
  • The seeing is 0.8-1.0 in the images

04:16 We loose the guiding star for wind gusting. The telescope can not remain on this target, the TO suggest to slew to another target to change the orientation.

04:22 Slew to telluric standard for Zamorani_11.

04: executeLUCIscript.sh M009zCgr32_tell.acq 0093-0096

04:45 executeLUCIscript.sh M009zCgr32_tell.spec 0097-0108

04:49 executeLUCIscript.sh M009zCgr32_tell.acq 0109-0116
  • During the first attempt of the MOS acquisition we are not able to check whether the stars are within the reference boxes. This is due both to clouds passing combine with slow refersmall reference boxes (2.0x2.0).

05:35 executeLUCIscript.sh M009zCgr32_tell.spec 0117-0131

08:25 Script aborted. Guide star lost due to the wind
  • Sky no more clear. We stay on LUCI

08:35 Slew to Lafranca_A6
  • Preset failed. Target in the same position of Zamorani

08:42 Slew to LaFranca_A9
  • Run STDA9_1_J.acq 0132-0135

08:55 Run telluric script STDA9_200_zJspec.spec 136-137

08:59 Run science acq A9_1_J.acq 0138-0141

09:15 Run science script A9_1_200_zJspec.spec 0142-0149

10.02 Run acquisition script STDA5_1_J.acq 0150-0152

10:12 Run telluric script STDA5_1_200_zJspec.spec 0153-0156

10:16 Run acquisition script A5_1_J.acq
  • Aborted. The problem was in the lack of PA and other parameters.

10:20 Run acquisition script A5_1_J_new.acq
  • Preset faild, the target is on the SW.

WE CAN NOT POINT SOURCES ON THE SOUTH-WEST
  • We have no more targets for LUCI, apart those orientated towards the forbidden region of the sky.

10.35 We ask the TO to SWITCH TO LBC
  • We run LBC in blue mode only

10:59 We are authorized to use LBC
  • We slew to Grazian_30A

11:08 dofpia,/first,/X2 on scientific field Grazian

11:19 Run grazian_GOODSN_UR_16.OB
  • lack of FOCUS in the image lbcb.20130219.111734 (even if the previous dofpia converged)

11:26 dofpia,/onlyblue,/X2 on scientific field Grazian

11:32 Run grazian_GOODSN_UR_16.OB
  • uncheck Right Arm on the Power Control
  • uncheck Red channel and mirror in the OB execution
  • the average seeing is 1.0

12:10 dofpia,/onlyblue,/X2 on scientific field Grazian

12:16 Run grazian_GOODSN_UR_17.OB
  • The average seeing is 1.1
  • Interrupted after 2 exposures: twilight

12:37 We slew to the standard SA109
  • dofpia,/onlyblue,/X2

12:44 Run SA109_Sr.ob

12:50 The TO closes the dome

M009zCgr32.acq -- EleanoraSani - 18 Feb 2013
Topic revision: r13 - 31 Mar 2013, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback